A gas is enclosed in a cylindrical vessel fitted with a frictionless piston. The gas is slowly heated for some time. During the process, 10J of heat is supplied and the piston is found to move out 10 cm. Find the increase in the internal energy of the gas. The area of cross-section of the cylinder = 4 cm2 and the atmospheric pressure = 100 kPa.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. dQ = 10 J dV = A × 10 cm^3 = 4 × 10 cm^3 = 40 × 10^–6 cm^3 dw = Pdv = 100 × 10^3 × 40 × 10^–6 = 4 cm^3 du = ? 10 = du + dw ⇒ 10 = du + 4 ⇒ du = 6 J.